Circular No. 8392

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dw IN UGC 11394 T. Puckett and J. Newton report the discovery of an apparent su-

pernova (mag 17.2) on an unfiltered CCD frame taken with the Puckett Observatory 0.35-m automated supernova patrol telescope in Ellijay, GA, on Aug. 19.19 UT (and confirmed by Puckett on frames taken on Aug. 20.17 with a 0.60-m reflector). The new object is located at α = 19^{h}03^{m}38^{s}.25, δ = +27^{o}36′54′′.5 (equinox 2000.0), which is 25′′.3 east and 31′′.3 north of the center of UGC 11394. SN 2004dw was not present on images taken by Puckett on 2003 Apr. 23 and four other nights back to 2000 (limiting mag ∼ 20.0).

SUPERNOVA 2004dx IN MCG +07-37-36 T. Boles, Coddenham, Suffolk, England, reports the discovery of an

apparent supernova (mag 18.1) on unfiltered CCD images taken on Aug. 19.953 and 21.950 UT with a 0.35-m reflector. The new object is located at α = 18^{h}18^{m}06^{s}.15, δ = +39^{o}49′24′′.9 (equinox 2000.0), which is ≈ 2′′.4 east and 6′′.3 south of the center of MCG +07-37-36. SN 2004dx is not present on Boles’ images from June 27 and May 18 (limiting mag 19.5), and it is not present on Digitized Sky Survey red (1992) or blue (1989) plates.

V1187 SCORPII D. K. Lynch, R. W. Russell, Aerospace Corporation; S. J. Bus and

W. Golisch, University of Hawaii; and T. Harrison, New Mexico State University, report further infrared spectroscopy of this nova (cf. IAUC 8386) using SpeX (range 0.8–2.5 microns) at the Infrared Telescope Facility on Aug. 13.3 UT, when the nova’s magnitude was K ∼ 6.8. The spectrum was dominated by H I emission lines with a Doppler width (FWHM) of ∼ 3000 km/s; the profiles showed a broad and a narrow (∼ 1000 km/s) component and were fairly symmetric. Also present were the Lyβ-fluoresced O I lines and weak He I line emission at 1.0830 and 2.0581 microns. There was no evidence of thermal emission from dust at longer wavelengths or of the CO first overtone at 2.3 microns. With a decline rate of 2 visual magnitudes in seven days (based on the AAVSO light curve), this rapidly fading nova should begin showing coronal lines in the next few weeks.

2004 August 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green